OpenStudy (anonymous):

please help A gym has 2 kg and 5 kg weights. There are 10 disks in all. The total weight of 2 kg disks and 5 kg disks is equal to 29 kg. Find the number of disks of each kind in the gym. a. 5 kg disk: 17; 2 kg disk: 7 b. 5 kg disk: 3; 2 kg disk: 7 c. 5 kg disk: 3; 2 kg disk: 13 d. 5 kg disk: 7; 2 kg disk: 3

OpenStudy (anonymous):

this cannot be that hard

OpenStudy (anonymous):

want to solve as a system of equations or some other way?

OpenStudy (anonymous):

im not good at word problems at all I tried it and failed

OpenStudy (anonymous):

ok lets do it a couple ways, make your teacher happy first of all, there are ten in all, and you have 4 choices

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

a. 5 kg disk: 17; 2 kg disk: 7 you gotta be kidding there are only 10 disks, not \(17+7=24\)!

OpenStudy (anonymous):

c. 5 kg disk: 3; 2 kg disk: 13 not likely since \(3+13=16\) so now we know it is either B or D

OpenStudy (anonymous):

b. 5 kg disk: 3; 2 kg disk: 7 ok at least this one has the right number of disks, namely 10 lets compute the weight and see if we get 29 kg

OpenStudy (anonymous):

3 5kg disks gives \(3\times 5=15\) kg and 7 2kg disks gives \(7\times 2=14\) kg and \(15+14=29\) so i guess we have to check no further

OpenStudy (anonymous):

now we didn't actually solve the problem we just checked which answer is right would you like to go ahead and solve it as if we did not have choices?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok we have a couple things we could do lets call the number of 5 kg disks a variable , say \(x\) then since the total number of disks is \(10\) then number of 2 kg disks is \(10-x\) (since they add to 10) then the total weight would be \[5x+2(10-x)\] which we know is \(29\) so we can solve \[5x+2(10-x)=29\] for \(x\)

OpenStudy (anonymous):

lets solve and see that we get \(x=3\) as predicted remove the parentheses carefully and get \[5x+20-2x=29\] combine like terms for \[3x+20=29\] subtract \(20\) \[3x=9\] and so \[x=3\] as we knew

OpenStudy (anonymous):

okay so it would be D